OpenStudy (anonymous):

Use unit multipliers to convert 27 cubic feet per minute to liters per hour

OpenStudy (anonymous):

1 hr. = 60 min 1 liter = 0.03531 cu. ft. Therefore 27 cu. ft.=27 (cu. ft.)/min *(60 min/1 hr.)*(1 liter/0.03531 cu. ft.) =[27*60/0.03531](liters/hr.) =45879.35429 liters/hr